What is the amount of potassium acetate to be added to 1 litre of N/10 acetic acid to make a buffer solution of pH=4 using acetic acid and potassium acetate? The dissociation constant of acetic acid =2*10^-5

Grade:11

1 Answers

vikas askiitian expert
13 years ago

let the amount of potassium acetate  = m grams

M(CH3COOK) = m/92 .........1

normality of acetic acid = molarity*1

M( CH3COOH) = 0.1               

Ka = [salt][H+]/[acid]

  PH =4 means [H+] = 10-4

Ka = M(CH3COOK][H+]/[CH3COOH]

2*10-5 = [m/92][10-4]/[10-1]

 m= 1.84grams